11) Suppose researchers obtained a random sample of 100 nurses, a random sample of 100 physicians, and a random sample of 100 other health-care providers. They asked the providers whether or not they agree with the statement, It is permissible to withdraw life- sustaining treatment without consent in certain circumstances. Next, the researchers performed a chi-square test to evaluate the association between the binary variable of agreement with the statement about end-of-life care and profession (nurse, physician, other providers) pothesis for the researchers chi- square test is equivalent to the following: The probability that nurses agree with the statement, the probability that physicians agree with the statement, and the probability that other providers agree with the statement are all the same. b) * point) What is the degrees of freedom for their chi-square test? Explain your answer.
Answer #1

a) The given statment is not true because the null hypothesis is variables are equal.

b) The degree of freedom Chi-Square test is,

df = n-1 =300 - 1 = 299

n = 100 +100+100 =300